OpenStudy (destinyyyy):

Divide and simplify to the form a+bi... Divide and express the result in standard form. 6i/9-8i

OpenStudy (destinyyyy):

4i/1+i

OpenStudy (destinyyyy):

If I remember correctly im suppose to times the top and bottom by 1-i

OpenStudy (welshfella):

multiply denominator and numerator by the conjugate of the denominator

OpenStudy (welshfella):

yes that is correct

OpenStudy (destinyyyy):

4i(1-i)/ (1+i)(1-i)

OpenStudy (welshfella):

right

OpenStudy (welshfella):

now the denominator will work out to a real number

OpenStudy (destinyyyy):

I dont remember if I cancel out like terms between the top and bottom or if I distribute the 4i

OpenStudy (destinyyyy):

4i+4/2

OpenStudy (welshfella):

first compute the denominator (1 + i)( 1 - i) = 1 + i - i -i^2 = ?

OpenStudy (destinyyyy):

1-i^2 ....

OpenStudy (welshfella):

and what is i^2 equal to ?

OpenStudy (destinyyyy):

Seriously? -1.

OpenStudy (destinyyyy):

2+2i

OpenStudy (welshfella):

yes im serious i^2 = -1 so 1 - i^2 = 1 - (-1) = 2

OpenStudy (destinyyyy):

I already said 2 minutes ago 4i+4/2 .. Im way ahead of you. I already finished

OpenStudy (welshfella):

oh sorry i missed that

OpenStudy (welshfella):

2 + 2i is the final answer

OpenStudy (destinyyyy):

Its fine. But all im asking for is refreshers.

OpenStudy (welshfella):

OK you are doing pretty well

OpenStudy (destinyyyy):

Um thanks

OpenStudy (welshfella):

very well

OpenStudy (destinyyyy):

Im working on a different one and im a bit stuck at the end... Divide and express the result in standard form. 6i/9-8i

OpenStudy (destinyyyy):

54i-48/145

OpenStudy (welshfella):

this cant be simplfiesdany further 54i 48 --- - -- 145 145

OpenStudy (welshfella):

- is the standard form

OpenStudy (destinyyyy):

I got help in another question... And what you said is incorrect.